# Unexpected but informative: What fixation-related potentials tell us about the processing of ambiguous program code

### v1

@@ 3 Discussion
The primary goal of this study was to apply an FRP approach – an established technique from psycholinguistics – to achieve a better understanding of the neurocognitive processes that occur when reading ambiguous program code (i.e., code that contains atoms of confusion) by means of exploring parallels between the processing of program code and natural language processing in an ecologically valid experimental setting. Therefore, we investigated whether atoms of confusion trigger electrophysiological responses, which turned out to be similar to those engaged by unexpected but plausible inputs during language comprehension. In particular, our FRP analysis revealed a late positivity with a frontal scalp topography. It onsets around after the first eye fixation in the area containing the atom of confusion and was significantly larger for ambiguous code compared to unambiguous code snippets. Notably, this component resembles other late frontal positivities found in natural language processing studies, in which the late frontal positivity is elicited by unexpected words, often but not necessarily in highly constraining contexts [43, 44, 45]. As outlined in Section 1, the late frontal positivity can be dissociated from the P600, a late ERP component, that is elicited by words that are perceived as highly implausible during online comprehension [25, 46].
@@
These ideas align well with a plausible mechanism for processing the atoms of confusion: Atoms of confusion are infrequent in code [2] and hence generally unexpected, but they are not implausible, or incorrect, as they reflect syntactically correct code patterns that can be unambiguously executed by the computer. However, when a comprehender encounters an atom of confusion, they are likely to be unsure at first about its effect on the program execution and variables they need to keep track of for the task (i.e., they are unsure about how they need to update their situation model), which is in line with observations that relate ambiguous code patterns with increased cognitive load [15]. As a result, comprehenders access their long-term memory to remember how exactly to interpret the atom of confusion. If this retrieval is successful, the comprehender can confidently update their situation model. In the case of code without atom, the situation model will also be updated, but there is no point of uncertainty followed by confidence on how to do that. This account also generates new predictions to be tested in future studies: We would expect that the late frontal positivity should only be present in trials where the comprehender could successfully resolve the ambiguity. It should hence not show up in situations in which the atom of confusion contains a syntactically invalid code element, for which we would expect a P600 with a posterior scalp topography [40]. Also, we do not expect a late frontal positivity to show up when the comprehender remains uncertain about the semantics of the atom, or when they would be given a different task that does not require updating the situation model.
@@
In summary, we presented, to our knowledge, the first-ever study that employs an FRP approach to explore the processing of ambiguous and potentially confusing elements in program code. The results suggest that the brain engages highly similar mechanisms for processing unexpected but plausible elements in program code and natural language comprehension. This implies that studies on program comprehension can rely on established ERP components in natural language processing. The FRP approach offers new research perspectives with high ecological validity, focusing on the temporal characteristics of online program code comprehension. Based on this common methodological ground, methods and findings from program comprehension research may be transferred back to research on natural language comprehension.
@@ 4.2 Experimental variables
The independent variable of our experiment is the snippet type, which has two conditions: either ambiguous or unambiguous. As dependent variables, we measured for each snippet the comprehension time, the answer correctness, the subjective difficulty rating, the eye movements via an eye tracker, and the brain activity via an EEG system (see Figure 2).
@@ 4.3 Stimuli
The code snippets in our study are based on the snippet pairs from the study of Langhout and Aniche [1], which we adjusted to reduce the fixations and confounding neurocognitive processes in the FRP analysis. In a pilot study, we rated these snippet pairs considering three criteria regarding the snippet structure, area of the atom of confusion, and behavioral response (cf. supplementary information, Appendix A). Specifically, we required that, within each snippet pair, the snippets have a comparable structure to reduce the influence of confounding factors. This was determined by the number and length of the code lines. Furthermore, to facilitate the selection of suitable fixations for FRP analysis, each snippet in a pair must contain a single clear area of interest that contains either the atom of confusion (for ambiguous snippets) or a clarified equivalent (for unambiguous snippets). Finally, we tested whether the snippets were of a comparative complexity and whether they were, in principle, understandable to students of our target population (familiarity with the syntax of Java) in a pilot study. To this end, we set a threshold of answer correctness or subjective difficulty rating as easy for either condition. Snippets that showed a clear violation of the expected behavior by less correctness or fewer easy ratings were excluded from the stimulus materials.
We used snippets ( snippet pairs), each pair consisting of one ambiguous and one unambiguous snippet, which are semantically equivalent and differ only in the presence of an atom of confusion. We presented only one condition of each snippet pair to each participant ( snippets), this way, avoiding learning effects. In total, each participant viewed an equal number of snippets for both conditions ( snippets each).
@@ 4.4 Procedure
After the experiment, we conducted a semi-structured interview, including questions on the participants’ views about the experiment and each snippet.
@@ 4.5.3 Eye-movement preprocessing
The eye-tracking data, specifically the fixations, are the basis for our FRP analysis to reliably detect processing of ambiguity at the correct moment. Missing a fixation or selecting an incorrect fixation would substantially affect the quality of our FRP analysis. Therefore, only trials with high data quality, including a data correction, were considered in our eye-movement analysis.
@@ 4.6.1 Behavioral analysis
We analyzed the behavioral data using linear mixed effects models. As a model family, we chose a Gaussian error model with logarithmic values for fixation duration. Answer correctness and subjective difficulty rating were coded as binary variables (, and , , respectively) and analyzed using binomial mixed effects models. For each behavioral measure, the regression models included the fixed parameter condition to identify whether a trial was ambiguous or unambiguous (Condition: , ), as well as the block number (BlockNo: –) and the trial number within a block (ItemOrder: –) as well as their interactions to account for potential learning and habituation effects. Linear mixed effects models included random intercepts and random slopes for condition under participants and items; the fixed effects structure was reduced using backward model selection based on the Akaike information criterion and a significance threshold of .
@@ 4.6.2 FRP analysis
Supplementary information
The supplementary information provides detailed information on the stimulus modification and selection, the ERP analysis including methods and results, and the complete overview of the results from the behavioral analysis.
Acknowledgements We would like to thank all the participants of the study, who made this research possible, as well as our research assistants Youssef Abdelsalam, Yannick Lehmen, and Marcellius William Suntoro, who helped with the technical setup.
Declarations
@@
Ethics approval and consent to participate
Our study was approved by the ethics committee of the Faculty of Human and Business Sciences at our host institution (Saarland University). The experiment was conducted according to the relevant guidelines and regulations provided by the ethical review board. All participants provided their informed consent and received a compensation of 12 euros per hour, calculated on a quarter-hourly basis.
@@
We provide all relevant data in line with open data principles under a CC-BY license, respecting our local privacy laws (GDPR). Specifically, the datasets collected during the experiment and generated for the analysis are long-term archived in the Zenodo repository Dataset for “Unexpected but informative: What fixation-related potentials tell us about the processing of ambiguous program code”, https://doi.org/10.5281/zenodo.14229849.
@@
Author contribution
Annabelle Bergum, Anna-Maria Maurer, Norman Peitek, Regine Bader, Axel Mecklinger, Janet Siegmund, and Sven Apel developed the concept and the general research idea. Annabelle Bergum, Anna-Maria Maurer, Norman Peitek, Regine Bader, Axel Mecklinger, and Sven Apel devised the study design and formulated the method. Anna-Maria Maurer conducted the study. Anna-Maria Maurer, Norman Peitek, and Vera Demberg performed the data analysis. All authors contributed to the interpretation, writing, editing and reviewing. Anna-Maria Maurer and Annabelle Bergum prepared the visualizations. Axel Mecklinger, Regine Bader, Sven Apel, Vera Demberg, and Janet Siegmund provided the funding and supervised the project.
Tables
| Trial statistics | GLMER results for condition = ambiguous | ||||
| Trial count | Unambiguous: | Estimate | Statistic | Probability | |
| Ambiguous: | / value | value | |||
| Comprehension time | Unambiguous: | ||||
| Ambiguous: | |||||
| Answer correctness | Unambiguous: | ||||
| Ambiguous: | |||||
| Subjective difficulty | Unambiguous: | ||||
| rating | Ambiguous: | ||||
[…]
| Programming Experience | # | % | Frequency of programming activity | # | % | ||
| 1 | Year | 1 | 4.1 | Multiple times a year | 2 | 8.3 | |
| 1–3 | Years | 7 | 29.2 | Multiple times a month | 5 | 20.8 | |
| 3–5 | Years | 7 | 29.2 | 1–2 times a week | 4 | 16.6 | |
| 5–10 | Years | 8 | 33.3 | 3–5 times a week | 9 | 37.5 | |
| 10 | Years | 1 | 4.1 | Every day | 4 | 16.6 | |

### v2

@@ 3 Discussion
The primary goal of this study was to apply an FRP approach – an established technique from psycholinguistics – to achieve a better understanding of the neurocognitive processes that occur when reading confusing program code (i.e., code that contains atoms of confusion) by means of exploring parallels between the processing of program code and natural language processing in an ecologically valid experimental setting. Therefore, we investigated whether atoms of confusion trigger electrophysiological responses, which turned out to be similar to those engaged by unexpected but plausible inputs during language comprehension. In particular, our FRP analysis revealed a late positivity with a frontal scalp topography. It onsets around after the first eye fixation in the area containing the atom of confusion and was significantly larger for confusing code compared to clean code snippets. Notably, this component resembles other late frontal positivities found in natural language processing studies, in which the late frontal positivity is elicited by unexpected words, often but not necessarily in highly constraining contexts [45, 46, 47]. As outlined in Section 1, the late frontal positivity can be dissociated from the P600, a late ERP component, that is elicited by words that are perceived as highly implausible during online comprehension [27, 48].
@@
These ideas align well with a plausible mechanism for processing the atoms of confusion: Atoms of confusion are infrequent in code [11] and hence generally unexpected, but they are not implausible, or incorrect, as they reflect syntactically correct code patterns that can be unambiguously executed by the computer. However, when a comprehender encounters an atom of confusion, they are likely to be unsure at first about its effect on the program execution and variables they need to keep track of for the task (i.e., they are unsure about how they need to update their situation model), which is in line with observations that relate confusing code patterns with increased cognitive load [16]. As a result, comprehenders access their long-term memory to remember how exactly to interpret the atom of confusion. If this retrieval is successful, the comprehender can confidently update their situation model. In the case of code without atom, the situation model will also be updated, but there is no point of uncertainty followed by confidence on how to do that. This account also generates new predictions to be tested in future studies: We would expect that the late frontal positivity should only be present in trials where the comprehender could successfully resolve the ambiguity. It should hence not show up in situations in which the atom of confusion contains a syntactically invalid code element, for which we would expect a P600 with a posterior scalp topography [42]. Also, we do not expect a late frontal positivity to show up when the comprehender remains uncertain about the semantics of the atom, or when they would be given a different task that does not require updating the situation model.
@@
In summary, we presented, to our knowledge, the first-ever study that employs an FRP approach to explore the processing of confusing elements in program code. The results suggest that the brain engages highly similar mechanisms for processing unexpected but plausible elements in program code and natural language comprehension. This implies that studies on program comprehension can rely on established ERP components in natural language processing. The FRP approach offers new research perspectives with high ecological validity, focusing on the temporal characteristics of online program code comprehension. These results are significant for software engineering research, as they provide a foundation for a better understanding of code comprehension. Based on this common methodological ground, methods and findings from program comprehension research may be transferred back to research on natural language comprehension. The findings of our study, in particular, and this line of research, in general, also have far-reaching implications for the practice of software engineering. By shedding light on how programmers’ brains process and respond to confusing code, we can enable programmers to develop and maintain more secure and reliable software. Specifically, our line of research supports programming practice by providing empirical evidence for coding conventions that avoid confusion-inducing code patterns. Moreover, programming language designers can leverage these insights to design consistent language constructs that minimize the potential for confusion and resulting severe mistakes. Finally, educators can incorporate these findings into their teaching practices, ultimately educating future programmers to apply to more effective software development practices.
@@ 4.2 Experimental variables
The independent variable of our experiment is the snippet type, which has two conditions: either confusing or clean. As dependent variables, we measured for each snippet the comprehension time, the answer correctness, the subjective difficulty rating, the eye movements via an eye tracker, and the brain activity via an EEG system (see Figure 2).
@@ 4.3 Stimuli
The code snippets in our study are based on the snippet pairs from the study of Langhout and Aniche [13], which we adjusted to reduce the fixations and confounding neurocognitive processes in the FRP analysis. In a pilot study, we rated these snippet pairs considering three criteria regarding the snippet structure, area of the atom of confusion, and behavioral response (cf. supplementary information, Appendix A). Specifically, we required that, within each snippet pair, the snippets have a comparable structure to reduce the influence of confounding factors. This was determined by the number and length of the code lines. Furthermore, to facilitate the selection of suitable fixations for FRP analysis, each snippet in a pair must contain a single clear area of interest that contains either the atom of confusion (for confusing snippets) or a clarified equivalent (for clean snippets). Finally, we tested whether the snippets were of a comparative complexity and whether they were, in principle, understandable to students of our target population (familiarity with the syntax of Java) in a pilot study. To this end, we set a threshold of answer correctness or subjective difficulty rating as easy for either condition. Snippets that showed a clear violation of the expected behavior by less correctness or fewer easy ratings were excluded from the stimulus materials.
We used snippets ( snippet pairs), each pair consisting of one confusing and one clean snippet, which are semantically equivalent and differ only in the presence of an atom of confusion. We presented only one condition of each snippet pair to each participant ( snippets), this way, avoiding learning effects. In total, each participant viewed an equal number of snippets for both conditions ( snippets each).
@@ 4.4 Procedure
After the experiment, we conducted a semi-structured interview, including questions on the participants’ views about the experiment and each snippet. Finally, we paid the compensation of 12 euros per hour, calculated on a quarter-hourly basis.
@@ 4.5.3 Eye-movement preprocessing
The eye-tracking data, specifically the fixations, are the basis for our FRP analysis to reliably detect potential confusion at the correct moment. Missing a fixation or selecting an incorrect fixation would substantially affect the quality of our FRP analysis. Therefore, only trials with high data quality, including a data correction, were considered in our eye-movement analysis.
@@ 4.6.1 Behavioral analysis
We analyzed the behavioral data using linear mixed effects models. As a model family, we chose a Gaussian error model with logarithmic values for fixation duration. Answer correctness and subjective difficulty rating were coded as binary variables (, and , , respectively) and analyzed using binomial mixed effects models. For each behavioral measure, the regression models included the fixed parameter condition to identify whether a trial was confusing or clean (Condition: , ), as well as the block number (BlockNo: –) and the trial number within a block (ItemOrder: –) as well as their interactions to account for potential learning and habituation effects. Linear mixed effects models included random intercepts and random slopes for condition under participants and items; the fixed effects structure was reduced using backward model selection based on the Akaike information criterion and a significance threshold of .
@@ 4.6.2 FRP analysis
Acknowledgements
We would like to thank all the participants of the study, who made this research possible, as well as our research assistants Youssef Abdelsalam, Yannick Lehmen, and Marcellius William Suntoro, who helped with the technical setup.
@@ Acknowledgements
We provide all relevant data in line with open data principles under a CC-BY license, respecting our local privacy laws (GDPR). Specifically, the datasets collected during the experiment and generated for the analysis are long-term archived in the Zenodo repository Dataset for “Unexpected but informative: What fixation-related potentials tell us about the processing of confusing program code”, https://doi.org/10.5281/zenodo.14229849.